Machinery was acquired at the cost of $500,000. It has a useful life of 10 years and a salvage value of $50,000. Calculate the amount of depreciation of each of the first three years using the straight line depreciation method and declining balance method
Answer #1

Solution: Straight line deprecation method

Depreciation for first year

Depreciation =(cost- salvage value) /useful life

=($500,000 - $50, 000) ×1/10×12/12

=$45,000

Depreciation for second year

=($500,000 -$50,000) ×1/10×12/12

=$45,000

Depreciation for third year

=($500,000 - $50,000) ×1/10×12/12

=$45,000

Declining balance method

Useful life = 10 years -> straight line

Depreciation rate = 1/10 = 10% per year

Depreciation rate for declining balance method

=10% ×100%= 10% per year

Depreciation for first year

=$500,000×10% ×12/12=$50,000

Depreciation for second year

=($500,000 - $50,000) ×10%×12/12 =$45,000

Depreciation for third year

=($450,000 -$45,000) ×10%×12/12 =$40,500

Declining balance method

Book value at the beginning Depreciation rate Depreciation Expense Book value at the year end
(For first year) $500,000 10% $50,000 $450,000
(For second year) $450,000 10% $45,000 $405,000

(For third year)

$405,000

10% $40,500 $364,500